Personal tools
You are here: Home / HARPS-N Project / System AIV at La Palma

System AIV at La Palma

First HARPS-N light
 
First HARPS-N light
First stellar light of HARPS-N. The exposure was taken in engineering mode.
Daily meeting notes of the AIV phase of HARPS-N at La Palma are reported




















 


Thursday 5th of April

by Francesco Pepe

During the past three night the weather conditions have been very bad and no observations or test could be done. The bad news is that the guiding, centering and ADC could not be tested further. The good news is that the observation chain is almost complete, from the NSTS to the data reduction (see below a FP frame for fun). Many bugs still remain and need to be solved during the second commissioning, which will start on April 12th at night. For this occasion, the spectrograph will be put to vacuum operations on April 11th.


The remaining Team has left the ORM on Thursday 5th at 11 o' clock. The instrument was switched off and the detector warmed up for pumping.

See you on April 12th again and have happy Easter days!

Tuesday 3rd of April

by Francesco Pepe

The past days were quite 'chaotic' in terms of activity to try to converge on various aspects simultaneously. Instead of reporting chronologically, we give a status update of the various subsystems.

Daytime

FEU:
  • On spectral flat-field calibrations we discovered modulations of the continuum of almost 30% of the flux and periodicity of 65 pixels. Since they were not present in Geneva, we deduced that it had to be produced by parallel plates interference in the FEU light path of the calibration light. One component in the parallel beam is the ND-wheel. After removing the ND-wheels, the most of these modulations disappeared. A residual pattern of about the same periodicity but 5% modulation was still visible. we deduced that it must be due to the glass plate holding the pupil mask in the calibration-light path. Today, the FEU was modified to remove the mask and change the location of the ND-wheel in the convergent light beam. After installation of the modified calibration-light system, the interferences have completely disappeared. 
  • The fibers from the calibration unit to the FEU were inverted and caused the wrong fiber to be illuminated. We exchanged the fibers and corrected the labels.
 
Spectrograph:
  • Exposure meter was installed and aligned. Values are of the order of 100'000 cps for the ThAr (140'000 cps for fiber A and 80'000 cps for fiber B). The difference is probably due to the fact that there is not ND-filter on fiber A and to some efficiency of the spectrograph fibers themselves. Some cross-talk of about 1% between the channels was observed.
  • The spectrograph was prepared for vacuum and is now to be closed and evacuated. This will be done on April 11th, however, in order to avoid interference with on-going tests, since the focus will have to be adapted when going in vacuum.
  • Investigating the diffusion effect on the CCD in the red orders: The suspicion that the fiber cladding of the octagonal fibers may transmit light could be discarded by making tests using lab fibers (exactly the same results). The fact that we have cross-talk on the exposure meter of 1% (HARPS-S 0.1%) may indicate to the possibility that the 'diffusion'  comes from the parabola or the multilayer coating. This possibility and possible solutions are being investigated. Another cause could be the gradient coating on the chip.
Detector:
  • Naidu and Charles have tested and characterized the CCD. The programmables gains were set to 1, 1.7, 2.66 and 3.33. The measured gains for the two available modes have been measured to been 1.73 and 1.0 for the 'fast' and the 'slow' read-out modes, respectively. The measured RON is of about 4.5 e- and 2.5 e-, achieved after some shielding optimization.
Software:
  • By today, the sequencer is running and delivering the correct and complete fits headers.
  • Christophe and Nicolas are debugging the templates to make all of them pass and produce the correct frames.
  • Basic functionalities of the ICS (LCU and AG) are running correctly. Some stability issues noticed. Other functionalities are implemented but have not yet been tested.
  • Pointing of the telescope by the sequencer working correctly and precisely. Note that the pointing of the telescope can be as off as 30 arc sec if not 'calibrated'.
  • Through-focus template needs to be developed.
  • AG guider guiding correctly, but the table defining filter as a function of magnitude needs to be reviewed.
  • ADC not yet tested
  • Exposure meter and monitoring functions not yet tested

Nighttime

  • Communication with the telescope tested and working:
  1. Pointing
  2. Offloading
  3. ADC telemetry
  • Guiding tested on several magnitudes. Seems to work well up to 15th magnitude (when clear sky ;-)
  • Automatic hole detection not yet reliable and not yet optimized
  • Offloading tested but not yet optimized (rotation angle of the camera to be verified in the corresponding vi)

Friday 30 March

by Christophe Lovis

Daytime

Arrival of Francesco, Michel, Nicolas. Departure of Gao, Naidu, Andy B. Organization of the work for the week to come.
 
Control software:
  • Demonstration of the LCU and AG interfaces to TNG staff.
  • Acquisition of calibration frames and checks of the behaviour of the FEU and CU.

 

Thursday 29 March

by Christophe Lovis

Daytime

Spectrograph/detector:
  • CCD tests and characterization continued. Grounding further improved, vacuum pump identified as a significant noise contributor, especially at low readout speed. It will have to be unplugged during normal calibration and science activities.
  • Measurement of readout noise, gain, photon transfer curve and linearity for slow readout mode. Results are nominal.
Control software:
  • Communication tests between telescope, sequencer (centering) and autoguider. No more conflicts.
  • Further work on the creation of FITS frames with correct filename and hierarchical FITS header gathering information from the various subsystems. To be continued.
  • Update of user instructions and demonstration of the LCU and AG interfaces. Gao leaving on March 30.

 Nighttime

  • Acquisition and guiding tests performed during the first half of the night.
  • Further debugging of calibration templates. Some issues with the FEU/CU configuration identified: calibration neutral density filters still to be properly adjusted, selection and on-time of ThAr lamps to be checked, "dark" CU carriage position still sending Fabry-Perot light...
  • Further guiding and offload tests performed. Good long-term stability was achieved on a V=8.6 star. Tip-tilt mirror position sometimes driven to large values, but mostly stays close to the middle. Further checks of the offload commands needed. Integrated guiding image looks somewhat asymmetric, centroiding algorithm to be checked.


Wednesday 28 March

by Christophe Lovis

Daytime

Spectrograph/detector:
  • CCD tests and characterization started. Careful grounding of all components.
  • Measurement of readout noise, gain, photon transfer curve and linearity for fast readout mode. Results are nominal.
  • Examination of LED frame shows that something is masking one of the LEDs. Visual inspection (opening a side flange on the vacuum vessel) has revealed that the culprit is the "giraffe" that was left over in front of the camera. To be removed when Michel returns.
Control software:
  • Implementation of a target centering procedure in the Sequencer. An external python module is called to compute the offsets between target and fiber hole and to send these offsets to the telescope.
  • Creation of FITS frames with correct filename and hierarchical FITS header gathering information from the various subsystems.
  • Update of the table relating stellar magnitude, neutral density filter position and fiber hole coordinates.

 Nighttime

  • Acquisition and guiding tests performed during the first half of the night.
  • Tests of the target centering procedure. After some initial debugging the offsets sent to the telescope are correct and the target can now be nicely centered on the fiber hole before starting the guiding. This is an essential step towards smooth operations of the instrument!
  • Some conflicts appeared between processes when sending pointing and offset commands to the telescope (several servers running simultaneously). To be solved with TNG staff today.
  • Further guiding and offload tests performed. Thanks to proper centering, guiding can now start properly, quickly finds the fiber hole and stabilizes. After some debugging the offset computation formula seems to send correct offloads to the telescope. The offloads correctly bring the tip-tilt mirror back towards "zero" position. Exact movements of the telescope to be further checked, but overall stability and robustness are very encouraging.

 

Tuesday 27 March

by Christophe Lovis

Daytime

Spectrograph:
  • Final connexions and adjustments (e.g. LEDs available for flat-fielding)
  • Closing of the vacuum vessel (no vacuum operation yet)
  • Spectrograph fully available for tests and characterization
  • Cabling work in the detector/spectrograph cabinet
Control software:
  • Tests and debugging of the execution sequence for templates tun_A, tun_AB, obj_A and obj_AB. The automatic configuration of the FEU and CU seems to be working well. The neutral density filter position for the calibration fibers should be recalibrated. Further tests to be done with the other templates.
  • Progress on the assembly of the final FITS frames with complete FITS header gathering information from the various subsystems.
  • Implementation of initialization and startup templates for the whole instrument. These can now be executed from the NSTS.

 Nighttime

  • Acquisition and guiding tests performed during the first half of the night.
  • The focus was on the computation of the guiding offsets and tests of the offload commands to the telescope. The sending of the offsets to the TCS is working properly. Checks are ongoing on the computed offsets to verify that the tip-tilt mirror position is properly taken into account in the computation. To be further tested tonight.
  • Tests of the ADC and neutral density filter positions are ongoing. The table relating stellar magnitude to neutral density position needs to be revised.


Monday 26 March

by Christophe Lovis

Daytime

Spectrograph:
  • Final cabling of pumping system
  • 3T window adaptation
  • Optical table removed from VV1 for final adjustments, gluing and screwing
  • Further adjustments and checks of detector vacuum
Control software:
  • Tests and debugging of the execution sequence for templates tun_AB, obj_A and obj_AB. The automatic configuration of the FEU and CU seems to be working well. Further tests to be done.
  • Acquisition procedure now working as foreseen (Sequencer receives acquisition image and the user clicks on the target).

 Nighttime

  • Guiding tests performed during the first half of the night.
  • The main difficulty is the lack of a centering procedure once the target has been selected. At the moment guiding can only start if the target is within a 100x100 (~17x17 arcsec) pixel window around the fiber hole. For some unknown reason the pointing was repeatedly off by ~1 arcmin and there is presently no way to automatically compute and send offsets to the telescope to properly center the target. Centering the target with the Dolores acquisition GUI does also center it properly for HARPS-N. The telescope pointing model will be checked. The implementation of a centering procedure is mandatory.
  • Tests of the ADC and neutral density filter positions have started. To be continued. Proper configuration of the folding mirror in the FEU for all acquisition templates should also be checked.

 

Sunday 25 March

by Christophe Lovis

Daytime

  • Various activities around the spectrograph
  • CCD tilt further adjusted to optimize focus in all parts of the spectrum
  • Problem of light "diffusion" around some of the spectral orders further investigated. The problem always affects the same orders, all on the red part of the chip, and is present whatever the light source (tungsten, ThAr, FP). It does not seem to depend on flux level (no diffusion-flux correlation when varying flux level or from order to order). No periodic pattern is seen, affected orders seem to be distributed at random. The problem persists whatever readout mode is chosen (slow, fast, 1 channel, 2 channels). Investigations ongoing to determine whether the problem comes from the CCD or the optics.

 

Saturday 24th of March

by Francesco Pepe

Daytime

Spectrograph:
  • Detector cold and running
  • All detector subsystems connected (controller, shutters, etc.)
  • Spectrograph aligned:
  1. Blaze angle adjusted such that to the right and left of each order there is equal intensity
  2. Rotation of the detector around optical axis verified. Orders in the middle of CCD about parallel to border between two read-out zones. No adjustment was necessary.
  3. Fiber rotated such that the two fibers are vertical on the chip. The guiding screw in the exit had to be removed and the fiber rotated by about 5 degrees clock-wise with respect to vertical axis as seen from the parabola.
  4. The height of the detector was adjusted (150 µm lower) such that the reddest order is safely at 13 pixels from edge and no order is cut in the center of the CCD.
  5. Tip and tilt are adjusted to achieve best image quality all over the CCD.
  • Some strange diffusion effect detected on the CCD. The orders and the spectral lines seems to be broader than others at the basis of the line. Some orders in the red part of the CCD seem to be affected much more than the others. To be investigated.

 Nighttime

  • Guiding started and tested on an 8th magnitude star. First working, then a problem occurred and it was not possible to make a startguiding anymore. After some attempts, we came back to nominal situation.
  • Acquisition of a V = 6.5 mag start was done. Guiding was started and a spectrum recorded: FIRST LIGHT with HARPS-N! Beautiful spectrum of expected SNR: SNR = 156 in 2 minutes with seeing of 0.7 arcsec.
  • Spectrum with simultaneous Fabry-Pérot recorded
  • GJ1214 (V = 14.7 mag) acquired and guiding started. Guiding working well at about 10 Hz. After introduction of a delay of 500 ms guiding images looks even better and very stable!

 

Friday 23th of March

by Francesco Pepe

Daytime

Spectrograph:

  • Detector pressure to 5.7E-6. Leak tests confirms no leak.
  • Cryogenic system prepared for detector cooling
  • Cooling started at 19.30 pm. After some problems (LN2 line was not ok and had to be changed with spare), the cooling started correctly at the temperature of the CCD set to -110°C.
  • First frames were taken at -50°C, which confirmed the correct functioning of the CCD.

Nighttime

  • Guiding tests repeated, but small problems prevented to repeat the results of the night before. After some corrections, the system was handed over to Rosario, Christophe, Charles and Francesco. It was possible to point 4 different stars and to guide on the hole with excellent precision (see movie). A few aspects were noted, however:
  1. The tracking drift may be as high as 1 arc sec per 10 minutes, which requires the AG to be able to offload.
  2. The hole position seems to shift as a function of AG attenuator position, probably due to the parallel plates introduce by the ND-filter wheel. A look-up table for the hole position must be built.
  3. The operations are quite difficult without automatic centering of the telescope onto the fiber-head hole after acquisition.
  4. Other minor improvement requests have been noted and will be reported.
  • The Sequencer sequencer was tested to provide the TCS with the stellar coordinates. It is now possible to point the telescope from the NSTS through the Sequencer automatically.


Thursday 22th of March

by Francesco Pepe

Daytime

FEU:

  • Calibration fibers aligned
  • Fiber-head cleaned
  • Calibration fibers attached
  • FEU and CU closed

Spectrograph:

  • Detector head pre-tested against leaks. None detected
  • Pumping system prepared to evacuate detector using TurboCube
  • Pumping started for the night

Nighttime

  • Telescope pointed and M4 put in position.
  • AG tests done on an arbitrary reference pixel. Seems to work well if reference pixel within a few arc sec of the star- Otherwise the system goes to saturation.
  • Step function given by the telescope. Tip-tilt is following perfectly at gains 0.1 to 0.5. Steps up to 2 arc sec are corrected perfectly, 5 arc sec is too much.
  • Tests stopped and telescope given to TNG of Queue Observations


Wednesday 21th of March

by Francesco Pepe

General: Almost new moon and start of Spring ...

Daytime

FIRST LIGHT of Marta, Adriano's beautiful baby! A star was born. Welcome Marta in this world! The HARPS-N Team wishes you and your family all the best!

FEU:

  • Alignments verified but nothing changed

Spectrograph:

  • Detector head and cryostat installed (very critical) and pumped with leak detector. Look good for the moment.
  • Minor small works done on the spectrograph

Nighttime

  • Telescope pointed and M4 put in position. No star.
  • Problem detected on M4 which had to be reset manually
  • Star acquired and visible on guiding camera close to Fiber A position. FIRST LIGHT of AG.
  • Focus by 1 mm compared to Dolores. Star now focused, but double. An oscillation of the (active) Derotator was the cause. After opening the loop of the derotator and keeping it at constant moment, the problem was disappeared.
  • Communication problem between ICS and FEU cabinet prevented to do guiding tests. Later solved and star re-acquired.
  • AG tests postponed to day after.

 

Tuesday 20th of March

by Francesco Pepe

FEU:

  • In the afternoon the alignment was checked with the second laser
  • Fiber source was set up and image quality checked: ok (145 µm fiber-source image almost 100% into 150 µm fiber-head hole)
  • With three different colors through fiber source, the ADC offset position was checked: ok. Zenith position corresponds to 180° for both ADC prisms
  • Request to ICS: At initial setup the ADC should be put to 180°.
  • All FEU cables connected

Hardware:

  • Newport mechanism controller access problem solved
  • Network-to-serial converter port: Geneva's devices have no problem. However, the small converters on the FEU cabinet work need a long time to initialize themselves after changing the network address. Working now properly.

Software:

  • ICS: AG camera connected to ICS and working properly
  • Full-frame read-out mode of AG now working
  • WCS coordinates-producing script will be prepared by TNG. It shall add WCS coordinates to the acquisition frames. When working, the script shall be integrated into the AG software.
  • Sequencer-NSTS interface ready to be tested
  • Sequencer-TCS interface may be tested tomorrow afternoon, but is not a test goal for the first night on sky
  • NSTS: Catalog import tested and working
  • NSTS: User Manual being prepared

Spectrograph:

  • Lot of visits ;-)
  • Fabry-Prot cleaned and cables attached properly
  • Baffles for the flat mirror made
  • Detector being installed on the optical bench
  • Exposure meter finished and ready to be aligned

System:

  • 21st confirmed to be first-light night and test night on FEU
  • For 22nd another half night of tests will be requested to TNG. Further plan to be discussed with TNG

 

Monday 19th of March

by Francesco Pepe

General: Another excellent day ...

System:

  • Provisional list of alarms defined by Andy Born and to be reviewed by Gao and Francesco

FEU:

  • 4 mm shims put under the tip-tilt mirror to adjust pupil position in x-direction, which gave better centering of the fiber pupil on tip-tilt mirror.
  • L1 was not in face of the FEU reference position (z-direction). FEU at 205 mm and L1 at 200 mm. After discussion of possible solutions, L1 was moved to 205 mm (easiest solution). Then M4 was adjusted in position correspondingly and tilts to overlap pupils perfectly on the tip-tilt mirror and the tip-tilt mirror was tilted to center the two alignment lasers (parallel to optical axis of the telescope) on the fiber hole (field center).
  • Todo: Check image quality using the LED and ADC rotation zero position. As well, align calibration fibers.

Thermal Enclosure:

  • Logic air personnel arrived at La Palma
  • Air-conditioning installation started
  • Two TNG persons helping LogicAir

Spectrograph:

  • Fiber paths fixed on the wall and rooted inside 2T
  • Fabry-Pérot installed definitively
  • Exposure meter PMTs installed

Detector:

  • Baffles in front of the CCD exchanged (previous were vignetting by 0.5 mm)
  • Field lens rotated
  • Detector head cleaned and closed

Hardware:

  • Gb-link cables installed in the FEU cabinet
  • The two Lakeshore devices not responding, however

Software:

  • NFS, shared file system installed for autoguider, ICS, DRS
  • Nicolas Buchschacher arrived
  • Communication not working with Newport controllers (ports are visible but not readable)
  • NSTS executable installed locally. Christophe Lovis asked for the installation of the catalogs
  • DRS/NSTS/Sequencer machine named
  • Sequencer now running properly on this same machine
  • Todo for first FEU light on 21st of March (if possible)
  1.   Test communication Sequencer-NSTS and Sequencer-TCS
  2.   Make ADCs running correctly
  3.   'Publish' full AG frames and guiding frames on NFS for off-line analysis
  4.   Run the AG code and have the tip-tilt running
  5.   Possibly make Exposure meter running for these tests
  6.   Ask Emilio to use part of the nights on 21st, 22nd, if possible

Sunday 18th of March

by Francesco Pepe

General: Reduced activity during the weekend

Spectrograph:

  • Echelle grating fined tuned
  • Pumping system pipes installed definitively, also up to Fabry-Pérot
  • Camera lenses thermal compensator installed according to Fisba instructions at 20°C (measured temperature was 19°C)
  • Instrument aligned. Seems just perfect. We saw beautiful spectra, but were unable to use the camera correctly to take pictures of it ;-). In summary, we did:
  1. Baffles verified, some to be adjusted
  2. Position of intermediate spectrum on folding mirror centered within 1 mm (low)
  3. White pupil of echelle grating perfectly aligned on grism mask
  4. Position and size of spectrum in focal plane corresponds to theoretical values (lateral perfectly centered to better than 0.5 mm, vertically within 1 mm of position measured at Geneva, lateral size of spectrum 65 mm, focal point is 4 mm after reference surface - should be 5.5 mm)
  5. Echelle grating tilted by 1° more than nominal. The blaze angle seems to be 1° higher than nominal, when blaze maximum centered on the CCD (TBC)
  • Todo for spectrograph:
  1. Adjust baffle height in front of exit holder to + 6 mm
  2. Put small baffles above screws of folding-mirror holder
  3. Installa baffles on échelle grating support
  4. Install and align exposure meter
  5. Adapt exit holder screw to allow a range of 2 mm in addition (focus adjustment)
  • Detector prepared for installation of new masks at the NOT clean room


Saturday 17th of March

by Francesco Pepe

General: Reduced activity during the weekend

FEU:

  • Telescope pupil and fiber-head pupil provisionally aligned on tip-tilt mirror by shimming. On Monday the final shimming plate will be installed.

Spectrograph:

  • Parabola mounted on bench and aligned perpendicular to optical axis (given by laser referring to mechanical bench axis)
  • Echelle grating mounted on bench

 

Friday 16th of March

by Francesco Pepe

FEU:

  • Pupil alignment issue detected between fiber-head beam and telescope. Fiber-head beam (back illumination) not centered on the tip-tilt mirror (reference pupil). Christmas lights defining the telescope pupil not centered of the tip-tilt mirror (reference pupil). Two actions to be taken:
  1. Move the tip-tilt mirror along the entering optical axis to better center the fiber-head pupil on the tip-tilt mirror to avoid vignetting (since the tip-tilt mirror is located in the pupil, the fiber-head pupil does not need any additional clear aperture - all field positions are overlapping). The final goal is two first avoid vignetting and second to make the fiber-head pupil and the Christmas-light pupil overlap. A shimming of about 3 mm will be required
  2. Move M4 away along the entering optical axis to align the Christmas pupil with the fiber-head pupil. The lens mounting of L1 must be displaced correspondingly.
  • We have asked Adriano and the technician to come on Saturday (tomorrow) to help us doing this modifications.

Control hardware:

  • CCD-controller gains changed to 1, 2, 3, 4
  • Wiring of the SP and DU cabinets finished
  • Moxa PLC communication successfully tested
  • Charles leaving tomorrow

Software:

  • SDSU camera system on the network
  • GPS installed but not yet connected to the antenna. Address to be assigned
  • Network switches substituted. Ethernet cable inside the FEI cabinet to be changed since it does not support Gb-Ethernet
  • ICS installed and running. Port configuration for Lakeshore devices still to be done

System:

  • Andy Born preparing list of alarms and ranges. To be circulated soon

Spectrograph

  • Grism installed and aligned
  • Fabry-Pérot installed
  • Piping of the vacuum system going on.
  • Optical alignment tools prepared
  • Exposure meter prism glued

Thursday 15th of March

by Francesco Pepe

FEU:

  • L1 fitted in the derotator
  • Fiber source installed in the derotator
  • Fiber head illuminated from the fiber-exit end and watched into the guiding camera
  • Guiding camera focused on the fiber head using torch illuminating the fiber-head mirror
  • Derotator fiber source focused on the fiber head and used to adjust mechanically the tip tilt mirror (set to middle range) to center the light with the fiber-head hole
  • To be checked tomorrow: the pupil alignment (using the 'Christmas' light M1) and the ADC orientation

Fibers:

  • Fabry-Perot fibers pulled from CU to spectrograph

Control SW:

  • TCS-ICS communication tested for ADC and parallactic angle
  • Started to integrate into one software system
  • Network and hardware configuration working now
  • Netgear Switch still not working (compatibility with HP?). Will be replaced by a spare TNG HP switch.

Spectrograph:

  • Grism mounted in its support and cleaned
  • Fixed-wall of the 3T mechanically installed and electrically connected
  • Pumping system installed and pipes connected to a large extent
  • Preparing the Fabry-Perot location (to be installed above the ventilation channel) to the left of the vacuum vessel
  • Heating foils for the vacuum vessel feet glued in their supports

Hardware:

  • Static addressing fixed on the CU and SP cabinets

 

Wednesday 14th of March

by Francesco Pepe

General: Again a very good day. The HARPS-N Team is great, and the TNG support fantastic. Thanks to everybody!

FEU:

  • All cables between cabinet and FEU attached and mechanisms tested. Cables, mechanisms and controllers are all ok.
  • Fiber head brought to the FEU and connected to it. Fiber fed with laser light from behind. Alignment looks already great.
  • Tomorrow optical alignment to be completed

Control HW and SW:

  • Network now working and IP mapping going on
  • Requested to OG to switch all components to static IP
  • Rooting for the GPS checked. Longer cable to be procured by Carlos
  • TNG configured the network switches for the DU and SP cabinets
  • All temperature sensors of the spectrograph wired and connected to controllers. All fine.
  • N-TCS integration carrying on
  • Autoguider running for the optical alignments of the FEU; therefore it will be difficult to make remote tests tomorrow.

Fiber run:

  • Fiber run performance checked: efficiency, FRD, etc. Everything is as in Geneva.
  • Fiber pulled from 1T to the FEU. Fiber length is perfect, no issue.

Spectrograph:

  • Final alignment of the optical bench with respect to the detector and distance between detector head and cryostat adjusted
  • Mechanical assemblies mounted on the bench (lens mounted, parabola support, etc.)
  • Started preparing the boxes for shipping back some of the OG material
  • Small mirror for the exposure meter being prepared. Mask glued on the sides
  • Preparation of the grism support started

DRS and scientific Software:

  • Telecon to discuss archiving and installation concept for DRS and NSTS
  • ICS machine installed and ready to receive control software

 

Tuesday 13th of March

by Francesco Pepe
General:

  • Arrival of Andy Born, Andy Vick, Francesco Pepe in the early morning

FEU:

  • All tilt and y-z alignments done. All ok, apart from a decentering in z on the alignment target on the FEU (about 2.5 mm lateral misalignment). May be adjusted with M4, to be checked whether critical. Impact to be estimated by ray tracing.
  • Even without any internal alignment and the tip-tilt switched off, the reference beam falls at 1 mm from reference hole in dummy fiber head! 
  • Tomorrow moving and connecting electronics before final alignment
  • Derotation tested using a laser beam while pointing the telescope. Beam stayed absolute stable.
  • Ethernet cables pulled to the derotator and tomorrow sockets will be mounted by TNG at the FEU cabinet

Electronics hardware and software:

  • Bad switch exchanged. It is possible to remote log-in on the ICS machines now.
  • Parallactic angle directly provided by TNG system
  • Offset tests to be done to verify 'pointing model' (direction for off-loading as a function of pointing coordinates). Will save on sky time.

Fiber Run:

  • Alignment of the fiber head verified and still perfect, as in Geneva.
  • A rope was pulled through the building to pull back the fiber run. 

Vacuum System:

  • Mechanical axis of the bench aligned with respect to the vacuum vessel motion axis.

1T:

  • Heating system installed
  • Room cleaned
  • Wiring of the spectrograph cabinet done
 

Monday 12th of March

by Ian Hughes 

  • Alignment of the front end unit in rotation has been done. Rx and Rz are now ok.
  • The rotation of Ry still have to be corrected of 0,3°.
  • Facility to redefine fiber hose on guiding camera has been done.
  • Fibers for network are all connected except for the switch.
  • Electric installation is now completed in 2T and nearly finished in 1T.
  • 3T has been unloaded for the container and been placed in 2T at the right position to determine the correct position of the fixed wall. This for drilling the floor before cleaning.
  • 3T is now stored under the Gazebo. The fans installation had started.
  • The optical table has been placed in the vacuum vessel for a first alignment.
  • 2T and the Gazebo has been completely cleaned and we are now in "clean room" configuration.

Sunday 11th of March

by Ian Hughes

  • The fixed wall of 3T is in 2T and ready to be fixed.
  • The vacuum vessel has been moved in place and been adjusted regarding the THK rails.
  • The vacuum vessel has been put back at atmospheric pressure and opened.
  • Ones the vacuum vessel was opened, an other check between the inside rails and the THK rails has been done.
  • The optical table has been brought inside 2T, but still in the Climafol envelope to protect from dust. For that, the centre of the vacuum vessel has been removed outside 2T and brought back in. (the hoist has been used intensely today !)
  • The roof of 2T is now closed.
  • The cover of the vacuum vessel (VV3) is stored under the "Gazebo".
  • The electric installation in 2T and the network is finish. Thank you to all the TNG team for they great job and they support during the week-end !!!
  • The heating feet system is cabled.
 

Saturday the 10th of March

by Ian Hughes

  • The vacuum vessel has been unloaded from the container. CT2 is empty and will be picked up Friday when Logicair's container comes up.
  • The hoist has been installed and connected by the people of TNG.
  • The I-beam of the hoist has been cleaned.
  • In 2T, the plastics again the wall and over the rails have been removed, some silicon seals have been done all around at the bottom of the wall, the roof has been opened for the hoist and a rough cleaning has been done inside and on the roof.
  • The THK rail have been aligned with a good accuracy. On side have moved with the pouring of the concrete.
  • The motion screw of the vacuum vessel has been installed.
  • At the end of the day, the vacuum vessel has been placed in 2T, but not in definitive position.
  • The detector unit has been connected and test. No issues, the pictures are the same as in Geneva at 20°C.
  • The spare components of the controller have been tested. No issues.
 

Friday the 9th of March

by Ian Hughes 

Telescope side:

  • The focus of the auto-guiding camera has been checked.
  • The Front End Unit has been closed and all the checks have been done previously. No issues.
  • The Front End Unit has been placed from horizontal to vertical position and has been installed on the frame. No issues. (see pictures form David)
  • The rotation has been done to checked the clearance. No issues.
  • The mounting brackets have been removed.
  • Some measurements have been done at the end of the day. The XYZ position is in 1 or 2 mm. The tilt is in 0.7° for Rz and 1.4° for Rx.
  • David and Angus are back on Monday.

Spectrograph side:

  • OG team arrived at 10:30 am at TNG, just after the container CT2.
  • Nothing inside the container has moving during the transportation. Everything is in good condition.
  • The container CT2 is empty except for the vacuum vessel. (This will be done tomorrow)
  • The carriage of the hoist has been installed on the beam. (The hoist will be hanged tomorrow)
  • A few boxes has been unpacked and the equipment stored in different places.
  • The detector unit has been unpacked and placed on the table in the rest room, ready to be connected tomorrow.
  • The second and last coating of paint on the floor of 2T has been done by Carlos from TNG.

 

Files
Guiding movie   22.3 MB  
Short movie of the first guiding sequence on the fiber-head hole
Fabry-Perot raw image   33.7 MB  
Raw image of a Fabry-Pérot spectrum taken with HARPS-N. Both fibers have been illuminated with the FP.
« August 2021 »
August
MoTuWeThFrSaSu
1
2345678
9101112131415
16171819202122
23242526272829
3031